A prisim is completely filled with 2304 cubes that have edge length of 1/4 cm. What is the volume of the prisim?
1
Expert's answer
2020-02-21T09:49:30-0500

Let we have a prism which completely filled with 2304 cubes that have edge length of 1/4. Then volume of 1 cube is

Vc=a3=(1/4)3=1/64 cm3. The volume of prism is Vp=2304Vc=2304*1/64= 36 cm3.